OpenStudy (anonymous):

Sam wrote this expression to describe the number of apples he gave away to his friends. 12 – a Which situation could be described by this expression? A. Sam does not know how many apples he has. He knows he gave away 12 apples. B. Sam knows he has 12 apples left. He does not know how many apples he gave away. C. Sam knows he has 12 apples left. He gave away half of his apples. D. Sam knows he had 12 apples to start. He does not know how many apples he gave away.

OpenStudy (anonymous):

that is question 2 of 8

OpenStudy (anonymous):

3. (1 pt) Riva wrote this expression to describe the amount she paid for each movie ticket. 25 ÷ t Which situation could be described by this expression? A. Riva knows that each ticket cost $25. She does not know how many tickets she bought. B. Riva knows she paid a total of $25. She does not know how many tickets she bought. C. Riva knows she paid a total of $25. She does not know how much popcorn and a drink will cost. D. Riva knows she paid $25 a ticket. She does not know the total amount she spent.

OpenStudy (anonymous):

4. (1 pt) Percy wrote this equation to represent the total number of books he carried in two bags. 12 = 7 + b ⁢       Which situation could be described by this equation? A. Percy knows he has a total of 12 books. He knows he has 7 books in one bag. He does not know how many books he has in the other bag. B. Percy knows he has a total of 12 books. He knows he has 7 books in one bag. He does not know how many more books he has in the first bag compared to the other bag. C. Percy knows he has 12 books in one bag. He knows he has 7 books in the other bag. He does not know how many books he has altogether. D. Percy knows he has 12 books in one bag. He knows he has 7 fewer books in the other bag. He does not know how many books he has altogether.

OpenStudy (anonymous):

5. (1 pt) Jason wrote this expression to describe the total amount he will spend on a magazine and a book. s + 11 Which situation could be described by this expression? A. Jason doesn't know how much the magazine costs. He knows the book costs $11 more than the magazine. B. Jason doesn't know how much the magazine costs. He knows the book costs $11. C. Jason doesn't know how much the magazine costs. He knows he will spend a total of $11 on both the magazine and the book. D. Jason doesn't know how much the magazine costs. He knows he is only allowed to spend $11.

OpenStudy (anonymous):

6. (1 pt) Jefferson wrote this expression to describe the number of papers he still has to deliver. Which situation could be described by this expression? A. Jefferson knows he has delivered 45 papers. He does not know how many he has left to deliver. B. Jefferson knows he has to deliver 45 papers. He does not know how many he has already delivered. C. Jefferson doesn't know how many papers he has delivered. He knows he has 45 left to deliver. D. Jefferson doesn't know how many papers he has delivered. He knows it is 45 less papers than he delivered yesterday.

OpenStudy (anonymous):

Jayna wrote this expression to describe the number of pages she needs to read each day to finish her book. Which situation describes what p and 7 mean? A. Jayna does not know how many pages she needs to read. She know she has already read for 7 days. B. Jayna does not know how many pages she needs to read. She knows that she has 7 days to finish her book. C. Jayna does not know how many days she has to finish reading. She knows she has to read 7 more pages. D. Jayna does not know how many pages she has already read. She knows she has to read 7 more.

OpenStudy (anonymous):

8. (1 pt) Each person was asked to take 2 red tiles and 5 blue tiles to create a design. The equation below represents the total number of tiles taken by each person. Which equation would represent the total number of tiles taken by 8 people? A. t=8+2+5 B. t=8* (2+5) C. t=2+5*8 D. t=(8*2)+5

OpenStudy (anonymous):

D. should be the right one! Sam knows he had 12 apples to start. He does not know how many apples he gave away.